OpenStudy (anonymous):

can someone teach me how to turn 8/5 into 13/5 (this is a mixed number)

OpenStudy (anonymous):

add 1

OpenStudy (anonymous):

Always 1?

OpenStudy (anonymous):

just for this specific problem 8/5+x=13/5 x=(13/5)-(8/5) x=(13-8)/5 x=5/5 x=1

OpenStudy (anonymous):

what if I want put 42.2857 into a mixed number. Can you guide me please

OpenStudy (anonymous):

do u know how to write .2857 as a fraction?

OpenStudy (anonymous):

0.2857 = 2857/10000, right?

OpenStudy (anonymous):

422857/10000

OpenStudy (anonymous):

yes now first u would write the decimal as afraction like u did then add whatever is in front of it to that decimal so: 42+(2857/10000)

OpenStudy (anonymous):

next u want to make it a mix number so u need the 42 to have the same denominator as the fraction u just made.

OpenStudy (anonymous):

I dont get where did you get 42 from?

OpenStudy (anonymous):

its what's before the decimal point: 42.2857

OpenStudy (anonymous):

ahhh got it

OpenStudy (anonymous):

continue, please

OpenStudy (anonymous):

ok so now u want the 42 and the (2857/10000) to have the same denominator so we rewrite 42 as 42/1 so now its (42/1)+(2857/10000) now we multiply the top and bottom of the 42/1 by the denominator in the fraction we first made. the 10000 out of the 2857/10000 so its (42/1)(10000/10000)+(2857/10000)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

when we simplify that it becomes (420000/10000)+(2857/10000) now since the have the same denominator we can combine the fraction: (420000+2857)/(10000)

OpenStudy (anonymous):

and simplify: 422857/10000

OpenStudy (anonymous):

I think I got it. Gonna try some exercises and if I get stacked I ask you for help again ;)

OpenStudy (anonymous):

so thats the mathematical way to do it. if u have a decimal the easy way to do it is count how many numbers are behind the decimal point rewrite it without a decimal point then divide it by 1 and a zero for each number u counted: 42.2857 there's 4 numbers behind the decimal point so rewrite it without a decimal point: 422857 now divide that number by 1 and a zero for however many numbers are behind the decimal point. since there is 4 we divide by 10000. if it was 3 the we would divide by 1000. if it was 5 we would divide by 100000 and so on. since its 4 it becomes 422857/10000

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

I have to solve a problem with this kind of numbers and according to my book 42.2857 = 42 2/7

OpenStudy (anonymous):

what ur book wrote is not necessarily tru unless it told u to round off to the nearest hundred thousandth spot

OpenStudy (anonymous):

well I got 42 2/7 and I did the following 42*7 + 2 = 296/7 = 42.285714285714285714285714285714

OpenStudy (anonymous):

I think I messed thing up, sorry

OpenStudy (anonymous):

yes which islightlydifferent from 42.2857 because the 1 u solved keeps goin. what did the problem specifically ask u to do?

OpenStudy (anonymous):

sorry for not replaying faster. My dog unplugged the source cable of the computer. The problem says the following: I have a debt of $183, after one month I have payed off 42 2/7. What is the amount of money to pay left?

OpenStudy (anonymous):

oo then u have to do 183-(42+2/7) so it becomes: 183-42-2/7 now u can subtract the whole numbers first so it becomes: 141-2/7 so what we would dohereis take 1 from the 141: 140+1 now we can change that to have the same denominator asthe fraction 140+(7/7) so now we can subtract the 2/7 140+7/7-2/7= 140+(7-2)/7 140+5/7

OpenStudy (anonymous):

dollars

OpenStudy (anonymous):

Definitively, that is the answer. Thank you sooo much ;)